OpenStudy (anonymous):

find the following sec(-14pi/3)

OpenStudy (anonymous):

Are you looking for an angle, or exact value?

OpenStudy (anonymous):

the value like if its \[\frac{ \sqrt{3} }{ 2 }or \frac{ 1 }{ 2} \]

OpenStudy (anonymous):

positive or negative

OpenStudy (anonymous):

but im not really sure how to do this one because its sec

OpenStudy (anonymous):

sec is just 1/cos

OpenStudy (anonymous):

and that value is -1/2

OpenStudy (anonymous):

ok can you show it step by step please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Do you know how to convert radians to degrees?

OpenStudy (anonymous):

yes but i know that it lies in quadrant 3

OpenStudy (anonymous):

Yes, so if you would convert (-14pi/3) to degrees--> (-840 degree), you can find exactly what value it is.

OpenStudy (anonymous):

wait i jus did it and i got positive 2

OpenStudy (anonymous):

It turns out to be 240 degrees

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

since it's in the third quadrant, it must be negative.

OpenStudy (anonymous):

oh yea sin and cos are both negative in quadrant 3 right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok thank you so much